OpenStudy (anonymous):

Please help; will fan and medal factor 3x^2+9x

OpenStudy (dayakar):

write the 9 into prime factors

OpenStudy (anonymous):

I don't know what prime factors are

OpenStudy (dayakar):

the numbers which have only two factors are prime numbers for example 2 = 1*2 [ 1 and 2 are two factors] 3= 1*3 5= 1*5

OpenStudy (anonymous):

3*3

OpenStudy (dayakar):

6 = 1*6 = 2*3 [ 1,2,3,and 6 are factors]

OpenStudy (dayakar):

very good

OpenStudy (dayakar):

\[3x ^{2}+9x = 3 \times x \times x +3 \times 3 \times x\]

OpenStudy (anonymous):

?

OpenStudy (dayakar):

\[x ^{2} = x \times x\] is it clear

OpenStudy (anonymous):

Not really.. My assignment has asked for it to be factored like (_+_)(_-_)

OpenStudy (anonymous):

At least that's what it wanted for the last 14 questions...

OpenStudy (dayakar):

i explain this one

OpenStudy (dayakar):

3x^2 + 9x which has two terms 3x^2 and 9x

OpenStudy (dayakar):



OpenStudy (dayakar):

3x^2 +9x = (3x)x + 3*(3x) =(3x) (x+3) [ taking 3x common]

OpenStudy (anonymous):

@$oumya @SeanStewart16 Can either of you break this down more? I'm extremely confused.

OpenStudy ($oumya):

x^2=x*x

OpenStudy ($oumya):

3*x*x+9*x =3*x*(x+3)

OpenStudy (anonymous):

Will you medal daykar since I can only give one and they tried.